Solve the differential equation x^(3)p^(2)+x^(2)yp+a^(3)=0 and also obtain its singular solution if
1
Expert's answer
2020-04-14T15:21:06-0400

To begin with, we clarify that by expression "p" we mean the first derivative with respect to the variable "x" :



"p=\frac{dy}{dx}"

Now we rewrite the initial equation, expressing the variable "y" as a function of "x" and "p"


"x^3p^2+x^2yp+a^3=0\longrightarrow\\[0.3cm]
\left. x^2yp=-a^3-x^3p^2\right|\div\left(x^2p\right)\longrightarrow\\[0.3cm]
y=-\frac{a^3}{x^2p}-\frac{x^3p^2}{x^2p}\longrightarrow\\[0.3cm]\boxed{y=-\frac{a^3}{x^2p}-xp}"

We differentiate the resulting expression with respect to "x" , remembering that "p=dy\/dx" in order to exclude the variable "y" from the equation.



"\frac{d}{dx}\left|y=-\frac{a^3}{x^2p}-xp\right.\longrightarrow\\[0.3cm]
p=-p-x\cdot\frac{dp}{dx}-\frac{a^3}{x^3p}\cdot(-2)-\frac{a^3}{x^2p^2}\cdot\left(-\frac{dp}{dx}\right)\\[0.3cm]
p+p-\frac{2a^3}{x^3p}=-\left(x-\frac{a^3}{x^2p^2}\right)\cdot\frac{dp}{dx}\\[0.3cm]
2p\cdot\left(1-\frac{a^3}{x^3p^2}\right)=-x\cdot\left(1-\frac{a^3}{x^3p^2}\right)\cdot\frac{dp}{dx}\\[0.3cm]
2p=-x\cdot\displaystyle\frac{dp}{dx}"




"2p=-x\cdot\frac{dp}{dx}\longrightarrow\frac{dp}{p}=-\frac{dx}{2x}\longrightarrow\\[0.3cm]
\ln|p|=-2\cdot\ln|x|+\ln|C|\longrightarrow\\[0.3cm]
\ln|p|+\ln\left|x^2\right|=\ln|C|\longrightarrow px^2=C\\[0.3cm]
\boxed{p=\frac{C}{x^2}}"



Substitute the resulting ratio in the original equation



"x^3p^2+x^2yp+a^3=0\longrightarrow\\[0.3cm]
x^3\cdot\left(\frac{C}{x^2}\right)^2+x^2y\cdot\left(\frac{C}{x^2}\right)+a^3=0\longrightarrow\\[0.3cm]
\left.\frac{C^2}{x}+Cy+a^3=0\right|\times(x)\\[0.3cm]
\boxed{C^2+(xy)C+a^3x=0}"

which is a quadratic equation in "C" . Equating it's discriminant to 0, we get :



"D=x^2y^2-4a^3x=0\longrightarrow\\[0.3cm]
x\cdot\left(xy^2-4a^3\right)=0\longrightarrow
\left[\begin{array}{l}
x=0\\[0.3cm]
xy^2-4a^3=0
\end{array}\right."



ANSWER



"\left[\begin{array}{l}
x=0\\[0.3cm]
xy^2-4a^3=0
\end{array}\right."

are the singular solutions of the given differential equation.